OpenStudy (anonymous):

Bobby ran 3/10 of a mile in 1/1/2 mintues. How many minutes per mile is this?

OpenStudy (anonymous):

Thats not much, I can run 5/10 of a mile in only 7/8 of a minute.

OpenStudy (anonymous):

do you mean miles per minute or minutes per mile if it is the later, then do (3/2)/(3/10) = (3/2)*(10/3) = 5 minutes per mile.